# Analysis of the near-side ridge structure in pp collisions via Momentum-Kick Model

### v1

@@ Abstract
The near-side ridge structure has been observed in the long-range two-particle correlations in heavy-ion collisions, such as AuAu collisions at the Relativistic Heavy Ion Collider and PbPb collisions at the Large Hadron Collider (LHC). Hydrodynamic models have successfully explained the ridge structure in heavy-ion collisions, indicating the presence of Quark-Gluon Plasma (QGP). Interestingly, similar ridge structures have been detected in high-multiplicity proton-proton and proton-lead collisions, which are classified as small systems in the LHC experiments. Because small systems have been considered insufficient to generate QGP, the applicability of theories developed for heavy-ion collisions to small systems remains controversial. Assuming that kinematic effects play a more significant role in small systems, we expect that the Momentum-Kick Model (MKM) can provide a satisfactory explanation. This model elucidates the long-range and near-side ridge structure in dihadron $\Delta\eta-\Delta\phi$ correlation by explaining that jet particles kick and rearrange medium partons along the direction of the jets. In this study, we apply the MKM to explain high multiplicity proton-proton collisions at both 13 TeV and 7 TeV in the LHC over various ranges of momenta. Furthermore, we introduce multiplicity dependence in the model to account for the 13 TeV data at various multiplicity ranges. We conclude that the MKM effectively explains the near-side ridge structure observed in proton-proton collisions. The LHC has entered Run 3, achieving higher center-of-mass energies and better luminosity than Run 2. We offer $\Delta\phi$ correlation predictions for pp collisions at 14 TeV and suggest possible extensions of the MKM for future studies.
I INTRODUCTION
@@
The Momentum-Kick Model (MKM) is based on purely kinematic interactions between the near-side jets and medium partons [41, 42, 43, 44]. Given that jets consist of particles with exceptionally high transverse momentum, it is logical that these jet particles kick the surrounding medium partons and propel them along the direction of the jets. This process is similar to the collective motion observed in heavy-ion collisions. The MKM elucidates the ridge structure through this process. Since the MKM has been successfully applied to the experimental results in AuAu at 200 GeV [45], PbPb at 2.76 TeV [46], and high-multiplicity pp at 13 TeV [47], we need to scrutinize whether the model can effectively explain the ridge structure in small systems.
All LHC experiments are currently investigating the ridge structure in small systems across various environmental conditions, including multiplicity ranges, center-of-mass energies, $p_{T}$ ranges, etc. Thus, it should be verified whether the MKM can explain all the data with different environments. In this paper, we will apply this model to high-multiplicity events at 13 TeV and 7 TeV center-of-mass energies, as well as to various multiplicity ranges at 13 TeV, from all three major collaborations: A Large Ion Collider Experiment (ALICE), Compact Muon Solenoid (CMS), and A Toroidal LHC Apparatus (ATLAS).
We introduce the main concepts of MKM in Section II. In Section III, we employ the MKM to analyze high-multiplicity proton-proton collisions at 13 TeV, utilizing data from ALICE, CMS, and ATLAS altogether [48, 49, 50]. We also examine high-multiplicity pp collisions at 7 TeV from the CMS [49]. Moreover, Section III is divided into three Subsections: Subsection III.1 explains the data analysis in experiments, and Subsection III.2 presents the fitting results of high-multiplicity proton-proton collisions at 13 TeV and 7 TeV. In addition, we discuss proton-proton collisions at 13 TeV for various multiplicity ranges from the ATLAS experiment in Subsection III.3. Section IV explains our prediction for the dihadron correlation yield at a higher center-of-mass energy, 14 TeV.
II Momentum-Kick Model (MKM)
@@
The MKM is designed to explain the near-side ridge structure. It focuses on the effect of the near-side jet particles kicking neighboring medium partons. The collective motion is induced by the kicked medium partons resulting in the near-side ridge structure. Since small systems create a relatively small number of medium particles compared to heavy-ion systems, we expect that kinematic behavior is more dominant in pp or pPb collisions than in PbPb collisions. Consequently, we anticipate that the MKM can effectively describe proton-proton collisions.
In order to better observe the ridge component, we select the long-range $\Delta\phi$ correlation data of pp collisions, where the ridge component dominates over the jet contribution. The ridge component in the MKM is represented by the charged ridge particle momentum distribution per trigger jet:
$\left[\frac{1}{N_{\text{trig}}}\frac{dN_{\text{ch}}}{p_{T}dp_{T}d\Delta\eta d\Delta\phi}\right]_{\text{ridge}}$
$\qquad\qquad\qquad\qquad=\frac{2}{3}f_{R}\left\langle N_{k}\right\rangle\frac{dF}{p_{T}dp_{T}d\Delta\eta d\Delta\phi},$ (1)
@@
$\frac{dF}{p_{T}dp_{T}d\eta d\phi}$ $=\left[\frac{dF}{p_{Ti}dp_{Ti}dy_{i}d\phi_{i}}\frac{E}{E_{i}}\right]_{\mathbf{p}_{i}=\mathbf{p}-\mathbf{q}}$ (2)
×$\qquad\qquad\times\sqrt{1-\frac{m^{2}}{\left(m^{2}+p_{T}^{2}\right){\cosh}^{2}y}}.$
In this equation, $dF/p_{Ti}dp_{Ti}dy_{i}d\phi_{i}$ represents the normalized initial parton momentum distribution. The factor $E/E_{i}$ ensures Lorentz invariance and the conversion of rapidity to pseudorapidity is represented by $\sqrt{\left.1-m^{2}\middle/\left(m^{2}+p_{T}^{2}\right){\cosh}^{2}y\right.}$. Here, $\mathbf{p}_{i}$ denotes the initial parton momentum, and $\mathbf{p}$ implies the final parton momentum shifted by $\mathbf{q}$. The initial parton distribution represents the momentum distribution before collisions. As the beam particles travel in opposite directions with equal energy, the net momentum of particles in the beam direction is zero. We assume that the jet particles lie in the perpendicular direction of the beam axis, and the transverse momentum of the initial particles can be expressed as:
@@
where $A_{\text{ridge}}$ serves as the normalization constant and $T$ represents the temperature of the medium particles. Given that pions are expected to make up the majority of the parton composition, we set the medium parton mass ($m$) equal to the pion mass. The fall-off parameter ($a$) determines the slope of the decrease in the $1-x$ distribution. We examined the fall-off parameter with several values (0.05, 0.5, 1, and 5), and no significant differences were observed. Consequently, we set $a=0.5$, consistent with previous studies [46, 45], ensuring the general applicability of the MKM.
@@
$x=\frac{\sqrt{m^{2}+p_{Ti}^{2}}}{m}e^{\left|y_{i}\right|-y_{b}},$ (5)
where $y_{b}$ corresponds to the beam rapidity, defined as $y_{b}=\cosh^{-1}{\sqrt{s_{\text{NN}}}/2m_{N}}$, with $m_{N}$ representing the mass of beam particles. Since we aim to elucidate the ridge structure in proton-proton collisions, we set $m_{N}$ to the proton mass.
$f_{R}\langle N_{k}\rangle$, $q$, and $T$ are the major parameters in the MKM to explain the ridge structure in the dihadron $\Delta\eta-\Delta\phi$ correlation. Consequently, we will compare the parameter values with previous studies in Subsection III.2.
III ANALYSIS
III.1 Data Analysis in Experiments
@@
       ALICE [48]        CMS [49]        ATLAS [50]
[…]
       $p_{T}$ range        $41<p_{T}<$ 4        $40.1<p_{T}<$ 4        $50.5<p_{T}<$ 5
Table 1: The experimental conditions for ALICE, CMS, and ATLAS [48, 49, 50]. The high-multiplicity range of the pp collisions in the CMS is specified $N_{\text{trk}}^{\text{offline}}\geq 110$ for 13 TeV, and $N_{\text{trk}}^{\text{offline}}\geq 105$ for 7 TeV. In the case of ALICE, the high-multiplicity range at 13 TeV is selected based on the top $0.1\%$ of the total data. For ATLAS, the high-multiplicity range of pp collisions at 13 TeV is defined as $N_{\text{ch}}^{\text{rec}}\geq 90$ [50], and the data are available for different multiplicity ranges. Additionally, the ALICE and CMS provide data for several $p_{T}$ bins, whereas the ATLAS does in a whole set with no separate bin.
@@
III.2 Fitting results for high-multiplicity pp collisions at 13 and 7 TeV
Given that recent research on ridge structures in small systems has primarily focused on the high-multiplicity region, we begin our analysis by applying the MKM to high-multiplicity pp collisions at 13 TeV and 7 TeV. The major parameters of the MKM for describing the ridge structure in two-particle correlations are $q$, $T$, and $f_{R}\langle N_{k}\rangle$. The parameter $q$ represents the average momentum transfer from jet particles to medium partons per kick, and the temperature, $T$, denotes the medium temperature. $f_{R}$ and $\langle N_{k}\rangle$ are the ratio of surviving particles at the detector and the average number of collisions with medium particles, respectively, and they work as a single parameter such as $f_{R}\langle N_{k}\rangle$.
@@ III.2 Fitting results for high-multiplicity pp collisions at 13 and 7 TeV
First, regarding the temperature $T$, we can observe a significant difference even at the same center-of-mass energy of 13 TeV. We treated the temperature as a free parameter for the 13 TeV data from the ALICE and CMS, unlike the previous study that fixed it based on the $\langle p_{T}\rangle$ ratio from AuAu collisions at 200 GeV [47]. This adjustment was necessary because pp collisions at 13 TeV represent a small-system environment with significantly higher center-of-mass energy than AuAu at 200 GeV, 65 times higher energy, which makes the previous method less reliable. The abundance of data makes it possible to treat $T$ as a free parameter. Since this methodological difference also influences other parameters, a comparison between our results and those in Ref. [47] is not meaningful. Therefore, we will limit our comparison to the results of 13 TeV (3rd column) and 7 TeV (4th column). Although the temperature for 13 TeV was treated as a free parameter, the temperature for 7 TeV can be fixed using the $\langle p_{T}\rangle$ ratio:
@@
This product is close to 1, supporting their inverse proportionality. $q$ and multiplicity are anti-proportional, and center-of-mass energy and multiplicity are proportional to each other. Therefore $q$ and center-of-mass energy are anti-proportional. We will use this relationship when incorporating the multiplicity dependence into the model in Subsect III.3. However, additional research is required to understand the underlying reasons for this relationship fully, including in other collision systems.
@@
Figure 3 shows the MKM results fitted to the high-multiplicity pp collision data from the ALICE and CMS [48, 49, 50]. The solid lines represent the MKM results, and the symbols indicate the experimental data from ALICE and CMS. The circles and diamonds correspond to the 13 TeV pp collision data from ALICE (red) and CMS (black), respectively, while the squares show the 7 TeV pp collision data from CMS (grey). We can see that the MKM describes the experimental data well for the available $p_{T}$ ranges in both ALICE and CMS.
@@
III.3 Extension to Other Multiplicity Ranges

### v2

@@ Abstract
The near-side ridge structure has been observed in the long-range two-particle correlations in heavy-ion collisions, such as AuAu collisions at the Relativistic Heavy Ion Collider(RHIC) and PbPb collisions at the Large Hadron Collider (LHC). Hydrodynamic models have successfully explained the ridge structure in heavy-ion collisions, indicating the presence of Quark-Gluon Plasma (QGP). Interestingly, similar ridge structures have been detected in high-multiplicity proton-proton and proton-lead collisions, which are classified as small systems in the LHC experiments. Because small systems have been considered insufficient to generate QGP, the applicability of theories developed for heavy-ion collisions to small systems remains controversial. Assuming that kinematic effects play a more significant role in small systems, we expect that the Momentum-Kick Model (MKM) can provide a satisfactory explanation. This model elucidates the long-range and near-side ridge structure in dihadron $\Delta\eta-\Delta\phi$ correlation by explaining that jet particles kick and rearrange medium partons along the direction of the jets. In this study, we apply the MKM to explain high-multiplicity proton-proton collisions at both 13 TeV and 7 TeV in the LHC over various ranges of momenta. Furthermore, we introduce multiplicity dependence in the model to account for the 13 TeV data at various multiplicity ranges. We conclude that the MKM effectively explains the near-side ridge structure observed in proton-proton collisions. The LHC has entered Run 3, achieving higher center-of-mass energies and better luminosity than Run 2. We offer $\Delta\phi$ correlation predictions for pp collisions at 14 TeV and suggest possible extensions of the MKM for future studies.
keywords:
Momentum-Kick Model, Near-side ridge structure, High-multiplicity pp collisions, Two-particle correlations
PACS:
25.75.Gz, 25.75.Ld, 25.75.Nq
journal: Nuclear Physics A
[…]
[first] organization=Department of Physics, Inha University,addressline=100, Inha-ro, city=Incheon, postcode=22212, country=Republic of Korea
1 INTRODUCTION
@@
The Momentum-Kick Model (MKM) is based on purely kinematic interactions between the near-side jets and medium partons [50, 51, 52, 53]. Given that jet fragments consist of particles with exceptionally high transverse momentum, it is logical that these jet particles kick the surrounding medium partons, rearranging them along the direction of the jets. This process is similar to the collective motion observed in heavy-ion collisions. The MKM elucidates the ridge structure through this process. Since the MKM has been successfully applied to the experimental results in AuAu at 200 GeV [20], PbPb at 2.76 TeV [54], and high-multiplicity pp at 13 TeV [37], we need to scrutinize whether the model can effectively explain the ridge structure in small systems.
All LHC experiments are currently investigating the ridge structure in small systems across various environmental conditions, including multiplicity ranges, center-of-mass energies, $p_{T}$ ranges, etc. Thus, it should be verified whether the MKM can explain all the data in different environments. In this paper, we will apply this model to high-multiplicity events at 13 TeV and 7 TeV center-of-mass energies, as well as to various multiplicity ranges at 13 TeV, from all three major collaborations: A Large Ion Collider Experiment (ALICE), Compact Muon Solenoid (CMS), and A Toroidal LHC Apparatus (ATLAS).
We introduce the main concepts of MKM in Section 2. In Section 3, we employ the MKM to analyze high-multiplicity proton-proton collisions at 13 TeV, utilizing data from ALICE, CMS, and ATLAS altogether [5, 31, 3]. We also examine high-multiplicity pp collisions at 7 TeV from the CMS [31]. Moreover, Section 3 is divided into three Subsections: Subsection 3.1 explains the data analysis in experiments, and Subsection 3.2 presents the fitting results of high-multiplicity proton-proton collisions at 13 TeV and 7 TeV. In addition, we discuss proton-proton collisions at 13 TeV for various multiplicity ranges from the ATLAS experiment in Subsection 3.3. Section 4 explains our prediction for the dihadron correlation yield at a higher center-of-mass energy, 14 TeV.
2 Momentum-Kick Model (MKM)
@@
As schematically depicted in Figure 1, the Momentum-Kick Model (MKM) explains the near-side ridge structure by proposing that near-side jet fragments impart transverse momentum to nearby medium partons. These kicked partons then stream outward along the jet axis, form a collective motion, and hadronize into stable final-state particles. In high-energy collisions, pairs of high-energy trigger particles are produced in nearly opposite directions and then interact with the surrounding medium, producing cone-shaped jet fragments. Since away-side jet fragments undergo multiple scatterings, which makes modeling the individual kicks on the away side difficult; hence, the MKM mainly focuses on the near-side phenomena. Because small systems create a relatively small number of medium particles compared to heavy-ion systems, we expect that kinematic effects is more dominant in pp or pPb collisions than in PbPb collisions. Consequently, given that Ref. [54] has demonstrated the MKM’s success in describing PbPb results, we anticipate that it will be even more effective in describing pp collisions.
In experimental analyses, two particles are randomly selected to generate dihadron correlations. On the near side, three types of correlation components typically emerge. The first involves correlations among the near-side jet fragments themselves: since these fragments remain confined in a small cone, correlations predominantly appear at $\Delta\phi\approx 0$ and $\Delta\eta\approx 0$. This component can be largely excluded by using long-range data. The second component emerges between near-side jet fragments and near-side kicked medium partons: if the initial medium partons are distributed in a plateau shape within mid-rapidity, the near-side jet kicks them, resulting in a peak at $\Delta\phi\approx 0$ across a wide $\Delta\eta$ range. The third component appears among the near-side kicked medium partons themselves; these kicked partons are propelled outward in a cone with roughly twice the opening angle of the jet cone, which naturally results in a $\Delta\eta$ range nearly twice as broad. In addition, under the assumption that the initial medium partons form a plateau in mid-rapidity, the third component preserves the shape of the second component. Thus, the third component does not influence the behavior within mid-rapidity.
To focus on the ridge structure, we select long-range experimental data. Consequently, we will explain the measurements using the ridge component in the MKM, which represents the yield of medium partons kicked by the jet. The ridge component in the MKM is represented by the charged ridge particle momentum distribution per trigger jet:
$\left[\frac{1}{N_{\text{trig}}}\frac{dN_{\text{ch}}}{p_{T}dp_{T}d\Delta\eta d\Delta\phi}\right]_{\text{ridge}}=\frac{2}{3}f_{R}\left\langle N_{k}\right\rangle\frac{dF}{p_{T}dp_{T}d\eta d\phi},$ (1)
@@
$\frac{dF}{p_{T}dp_{T}d\eta d\phi}=\left[\frac{dF}{p_{Ti}dp_{Ti}dy_{i}d\phi_{i}}\frac{E}{E_{i}}\right]_{\mathbf{p}_{i}=\mathbf{p}-\mathbf{q}}\times\sqrt{1-\frac{m^{2}}{\left(m^{2}+p_{T}^{2}\right){\cosh}^{2}y}}.$ (2)
In this equation, $dF/p_{Ti}dp_{Ti}dy_{i}d\phi_{i}$ represents the normalized initial parton momentum distribution. The factor $E/E_{i}$ ensures Lorentz invariance, and the conversion of rapidity to pseudorapidity is represented by $\sqrt{\left.1-m^{2}\middle/\left(m^{2}+p_{T}^{2}\right){\cosh}^{2}y\right.}$. Here, $\mathbf{p}_{i}$ denotes the initial parton momentum, and $\mathbf{p}$ implies the final parton momentum shifted by $\mathbf{q}$. Although $\mathbf{q}$ can vary with the jet particle’s momentum, we treat it as an average constant parameter that represents the statistically averaged momentum kick. The initial parton distribution represents the momentum distribution before collisions. As two beams travel in opposite directions with equal energy, the net momentum of the two colliding particles in the beam direction is zero. We assume that the jet particles come out in the perpendicular direction of the beam axis, and that the transverse momentum of the initial particles can be expressed as:
@@
where $A_{\text{ridge}}$ serves as the normalization constant for the initial partons, and $T$ represents the temperature of the medium particles. Although the medium is composed of a variety of partons, we set the medium parton mass ($m$) to the pion mass because pions dominate the observed hadron composition. The fall-off parameter ($a$) determines the slope of the decrease in the $1-x$ distribution. We examined the fall-off parameter with several values (0.05, 0.5, 1, and 5), and no significant differences were observed. Consequently, we set $a=0.5$, consistent with previous studies [54, 20], ensuring the general applicability of the MKM.
@@
$x=\frac{\sqrt{m^{2}+p_{Ti}^{2}}}{m_{N}}e^{\left|y_{i}\right|-y_{b}}.$ (5)
An increase in the daughter particle’s rapidity $(y_{i}$ leads to an exponential rise in the value of $x$. This exponential behavior is determining the fraction of the parent’s longitudinal momentum that the daughter particle inherits during the production process [49]. In the Eq.(5), $y_{b}$ corresponds to the beam rapidity, defined as $y_{b}=\cosh^{-1}{\sqrt{s_{\text{NN}}}/2m_{N}}$, with $m_{N}$ representing the mass of beam particles. Since we aim to elucidate the ridge structure in proton-proton collisions, we set $m_{N}$ to the proton mass.
We investigate the jet-induced changes in the medium parton momentum distribution by comparing the initial and final partons. The final parton momentum distribution is obtained without $f_{R}$ and $\langle N_{k}\rangle$. Figure 3 represents how the distributions differ between the initial and final partons. Panel (a) illustrates the $p_{T}$ distribution in the long-range ($52<|\Delta y|<$ 5) and near-side ($|\Delta\phi|<1$) regions. We can observe that most medium partons initially have low transverse momentum. However, after being kicked by the jet fragments, their $p_{T}$ distribution is shifted upward by approximately $q$. Panel (b) presents the long-range $\Delta\phi$ distribution, revealing that medium partons, equally populated along $\phi$ initially, elongate toward the jet direction and converge near $\Delta\phi\sim 0$. Panel (c) focuses on the near-side $\Delta y$ distribution, indicating that partons with larger $\phi$ values are realigned toward smaller $\phi$, resulting in a significant increase in the final parton count in the near-side region. For clearer comparison, the initial curve is multiplied by $2.1$. Panel (d) displays the $\Delta y$ distribution over the full azimuthal range. A distinct clustering of medium partons emerges near $\Delta y\sim 0$, whereas the distribution away from $\Delta y\sim 0$ remains nearly unchanged. This behavior reflects that the jet pseudorapidity ($\eta_{\text{jet}}=0$) makes the jet-induced kicks along the $y$ direction relatively small. These results illustrate how jet-induced kicks can significantly reshape the transverse momentum, angular and rapidity distributions of medium partons.
Figure 3: Comparison of the initial and final momentum distributions. Panels (a), (b), and (c) show near-side distributions, while panel (d) covers the full azimuthal range. The plotted quantities are $p_{T}$ (a), $\Delta\phi$ (b), and $\Delta y$ (c, d). Black dashed curves correspond to initial partons, while red solid curves correspond to final partons. In panel (c), the initial distribution is multiplied by 2.1 for clarity.
$f_{R}\langle N_{k}\rangle$, $q$, and $T$ are the major parameters in the MKM to explain the ridge structure in the dihadron $\Delta\eta-\Delta\phi$ correlation. Consequently, we will compare the parameter values with previous studies in Subsection 3.2.
3 ANALYSIS
3.1 Data Analysis in Experiments
ALICE [5] CMS [31] ATLAS [3]
[…]
$p_{T}$ range $41<p_{T}<$ 4 $40.1<p_{T}<$ 4 $50.5<p_{T}<$ 5
Table 1: The experimental conditions for ALICE, CMS, and ATLAS [5, 31, 3]. The high-multiplicity range of the pp collisions in the CMS is specified $N_{\text{trk}}^{\text{offline}}\geq 110$ for 13 TeV, and $N_{\text{trk}}^{\text{offline}}\geq 105$ for 7 TeV. In the case of ALICE, the high-multiplicity range at 13 TeV is selected based on the top $0.1\%$ of the total data. For ATLAS, the high-multiplicity range of pp collisions at 13 TeV is defined as $N_{\text{ch}}^{\text{rec}}\geq 90$ [3], and the data are available for different multiplicity ranges. Additionally, the ALICE and CMS provide data for several $p_{T}$ bins, whereas the ATLAS groups all data into a whole set with no separate bins.
@@ 3.1 Data Analysis in Experiments
3.2 Fitting results for high-multiplicity pp collisions at 13 and 7 TeV
Given that recent research on ridge structures in small systems has primarily focused on the high-multiplicity region, we begin our analysis by applying the MKM to high-multiplicity pp collisions at 13 TeV and 7 TeV. The major parameters of the MKM for describing the ridge structure in two-particle correlations are $q$, $T$, and $f_{R}\langle N_{k}\rangle$. Jet particles scatter off medium partons several times while they pass through the medium. Whenever they scatter, they transfer the momentum to the medium partons. Even though the number of scatterings per jet particle is fewer in pp collisions than in AA collisions, jet particles transfer the momentum to the medium partons. $q$ is the average value of momentum transfer per kick over several scattering events and over the jet fragment particles. The temperature, $T$, denotes the temperature of medium partons, which is more like an average kinetic energy rather than a statistical one. $f_{R}$ and $\langle N_{k}\rangle$ are the ratio of surviving particles at the detector and the average number of collisions with medium particles, respectively, and they work as a single parameter such as $f_{R}\langle N_{k}\rangle$.
@@ 3.2 Fitting results for high-multiplicity pp collisions at 13 and 7 TeV
First, regarding the temperature $T$, we can observe a significant difference even at the same center-of-mass energy of 13 TeV. We treated the temperature as a free parameter for the 13 TeV data from the ALICE and CMS, unlike the previous study that fixed it based on the $\langle p_{T}\rangle$ ratio from AuAu collisions at 200 GeV [20, 37]. This adjustment was necessary because pp collisions at 13 TeV represent a small-system environment with significantly higher center-of-mass energy than AuAu at 200 GeV, 65 times higher energy, which makes the previous method less reliable. In addition, the abundance of data makes it possible to treat $T$ as a free parameter. The observed increase of $T$ in pp collisions relative to Au+Au can be attributed to the larger total momentum of the measured particles [37]. Although we similarly observe a rapid increase in $T$, the ratio $T/T_{\text{AuAu 200GeV}}$ falls from 3.08 to 2.38, reflecting the substantially different experimental environment compared to heavy-ion collisions [37]. Since this methodological difference also influences other parameters, a comparison between our results and those in Ref. [37] is not meaningful. Therefore, we will limit our comparison to 13 TeV (3rd column) and 7 TeV (4th column) results. Although the temperature for 13 TeV was treated as a free parameter, the temperature for 7 TeV can be fixed using the $\langle p_{T}\rangle$ ratio:
@@
This product is close to 1, supporting their inverse proportionality. $q$ and multiplicity are anti-proportional, and center-of-mass energy and multiplicity are proportional to each other. Therefore $q$ and center-of-mass energy are anti-proportional. We will use this relationship when incorporating the multiplicity dependence into the model in Subsect 3.3. However, additional research is required to understand the underlying reasons for this relationship fully, including in other collision systems.
@@
Figure 4 shows the MKM results for the $\Delta\phi$ correlation fitted to the high-multiplicity pp collision data from the ALICE [5] and CMS [31]. The solid lines represent the MKM results, and the symbols indicate the experimental data from ALICE and CMS. The circles and diamonds correspond to the 13 TeV pp collision data from ALICE (red) and CMS (black), respectively, while the squares show the 7 TeV pp collision data from CMS (grey). We can see that the MKM describes the experimental data well for the available $p_{T}$ ranges in both ALICE and CMS.
Figure 5: The MKM results for $Y^{\text{ridge}}$. Red squares and grey squares represent experimental data from ALICE pp at 13 TeV and CMS pp at 7 TeV, respectively. Dashed lines connect the corresponding theoretical results. The $p_{T}$ bins used for the theoretical results are chosen to match those of the experimental data.
Figure 5 presents the MKM results for the yield. Grey circles and red squares represent experimental data from CMS pp at 7 TeV [31] and ALICE pp at 13 TeV [5], respectively. Theoretical results are connected by dashed lines.
@@
3.3 Extension to Other Multiplicity Ranges